suppose the growth rate of production in country A follows Harrod-Domar Model.The capital K in country A is 60 in year 1.The incremental capital-output ratio v= 5 and the constant savings rates s=0.1. For country A,What is the amount of aggregate production in year 1?
Answer #1

As growth rate of Output

g = s/c

c = incremental K-Y ratio

g = .1/5 = .02

So output growth rate = 2%

Then, initially K/Y = 5

Y = 60/5 = 12

So as, .02 = (Y1-Y0)/Y0

.02 = Y1/12 - 1

1.02 = Y1/12

Y1 = 12.24